The Foundational 15 [LO2-1, LO2-2, LO2-3, LO2-4]

[The following information applies to the questions displayed below.]

Sweeten Company had no jobs in progress at the beginning of March and no beginning inventories. The company has two manufacturing departments—Molding and Fabrication. It started, completed, and sold only two jobs during March—Job P and Job Q. The following additional information is available for the company as a whole and for Jobs P and Q (all data and questions relate to the month of March):

Molding Fabrication Total
Estimated total machine-hours used 2,500 1,500 4,000
Estimated total fixed manufacturing overhead $ 10,000 $ 15,000 $ 25,000
Estimated variable manufacturing overhead per machine-hour $ 1.40 $ 2.20
Job P Job Q
Direct materials $ 13,000 $ 8,000
Direct labor cost $ 21,000 $ 7,500
Actual machine-hours used:
Molding 1,700 800
Fabrication 600 900
Total 2,300 1,700

Sweeten Company had no underapplied or overapplied manufacturing overhead costs during the month.

Required:

For questions 1-8, assume that Sweeten Company uses a plantwide predetermined overhead rate with machine-hours as the allocation base. For questions 9-15, assume that the company uses departmental predetermined overhead rates with machine-hours as the allocation base in both departments.

1. What was the company’s plantwide predetermined overhead rate?

2. How much manufacturing overhead was applied to Job P and how much was applied to Job Q?

3. What was the total manufacturing cost assigned to Job P?

4. If Job P included 20 units, what was its unit product cost?

5. What was the total manufacturing cost assigned to Job Q?

6. If Job Q included 30 units, what was its unit product cost?

7. Assume that Sweeten Company used cost-plus pricing (and a markup percentage of 80% of total manufacturing cost) to establish selling prices for all of its jobs. What selling price would the company have established for Jobs P and Q? What are the selling prices for both jobs when stated on a per unit basis assuming 20 units were produced for Job P and 30 units were produced for Job Q?

8. What was Sweeten Company’s cost of goods sold for March?
Answer #1

Requirement 1:

Plantwide predetermined overhead rate = Total Manufacturing overhead cost / Total Machine hours

Total Manufacturing Overhead cost

Variable manufacturing overhead cost = (2,500 machine hours * $1.40) + (1,500 machine hours * $2.20) = $6,800

Fixed manufacturing overhead cost = $25,000

Total manufacturing overhead cost = $6,800 + $25,000 = $31,800

Plantwide predetermined overhead rate = $31,800 / 4,000 machine hours = $7.95 per machine hour.

Requirement 2:

Manufacturing Overhead applied to Job P = 2,300 Machine hours * $7.95 per machine hour = $18,285

Manufacturing Overhead applied to Job Q = 1,700 machine hours * $7.95 per machine hour = $13,515

Requirement 3:

Total Manufacturing Cost of Job P

Particulars $
Direct Materials 13,000
Direct Labor cost 21,000
Manufacturing Overhead cost 18,285
Total Manufacturing Costs 52,285

Requirement 4:

Unit Product cost of Job P = $52,285 / 20 Units = $2,614.25

Requirement 5:

Total Manufacturing Cost of Job Q

Particulars $
Direct Materials 8,000
Direct Labor cost 7,500
Manufacturing Overhead cost 13,515
Total Manufacturing Costs 29,015

Requirement 6:

Unit Product cost of Job Q = $29,015 / 30 Units = $967.17

Requirement 7:

Selling Price of Jobs

Particulars Job P Job Q
Total Manufacturing cost $52,285 $29,015
Add: Markup @ 80% $41,828 $23,212
Selling Price $94,113 52,227

Selling Price per unit of Job P = $94,113 / 20 Units = $4,705.65 per unit

Selling price per unit of Job Q = $52,227 / 30 Units = $1,740.90 per unit

Requirement 8:

Sweeten company's cost of goods sold = $52,285 + $29,015 = $81,300

All the best...